# Searching for coupled, hyperlight scalars across cosmic history

### v1

@@ Abstract
Cosmological scalar fields coupled to the Standard Model drive temporal variations in the fundamental constants that grow with redshift, positioning the early Universe as a powerful tool to study such models. We investigate the dynamics and phenomenology of coupled scalars from the early Universe to the present to consistently leverage the myriad searches for time-varying constants and the cosmological signatures of scalars’ gravitational effects. We compute the in-medium contribution from Standard Model particles to the scalar’s dynamics and identify only a limited range of couplings for which the scalar has an observable impact on the fundamental constants without either evolving before recombination or gravitating nonnegligibly. We then extend existing laboratory and astrophysical bounds to the hyperlight scalar regime. We present joint limits from the early and late Universe, specializing to hyperlight, quadratically coupled scalars that modulate the mass of the electron or the strength of electromagnetism and make up a subcomponent of the dark matter today. Our dedicated analysis of observations of the cosmic microwave background, baryon acoustic oscillations, and type Ia supernovae provides the most stringent constraints on quadratically coupled scalars with masses from $10^{-28.5}$ to $\sim 10^{-31}\leavevmode\nobreak\ \mathrm{eV}$, below which quasar absorption spectra yield stronger bounds. These results jointly limit hyperlight scalars that comprise a few percent of the current dark matter density to near- or subgravitational couplings to electrons or photons.
@@ I Introduction
Scalar fields have cosmological consequences as well—even simply through gravity if they are sufficiently abundant. A cosmological density of the new scalar may be produced in the very early Universe, for instance through the misalignment mechanism [29, 30, 31, 32, 33, 13, 34, 35, 36]. Depending on its initial conditions and interactions, the scalar makes up at least a subcomponent of the cosmological dark sector today as matter, radiation, or something more exotic. In general, scalar fields redshift with the expansion of the Universe and their observational effects are more pronounced earlier in cosmological history—more so if the leading dependence of fundamental constants on the scalar field amplitude is quadratic [23, 26, 37, 38] or higher rather than linear. In an era of precision cosmology, early Universe observations are particularly apt probes of dynamical scalar fields and may explore parameter space complementary to that from traditional particle physics tests.
@@
We use natural units in which $\hbar=c=1$ and define the reduced Planck mass $M_{\mathrm{pl}}\equiv 1/\sqrt{8\pi G}=2.435\times 10^{18}\leavevmode\nobreak\ \mathrm{GeV}$. We employ the Einstein summation convention for spacetime indices; repeated spatial indices (Latin characters) are implicitly summed regardless of their placement. We also use upright boldface to denote spatial vectors. Unless otherwise specified, we fix a homogeneous Friedmann-Lemaître-Robertson-Walker (FLRW) spacetime with metric
@@ II.1 Varying constants with new scalars
The assumed suppression of terms linear in $\varphi$ could be achieved with a $\mathbb{Z}^{\varphi}_{2}:\varphi\mapsto-\varphi$ symmetry.
@@
Depending on the assumed symmetries, scalars may interact with fermions and gauge bosons via operators other than those parameterized by Eq. 8. For instance, a shift-symmetric pseudovector coupling is characteristic of (pseudo)Nambu-Goldstone bosons such as axionlike particles [68, 69, 70, 71]. A pseudoscalar coupling is also allowed by symmetries and is equivalent to the pseudovector coupling alone only in processes whose Feynman diagrams have a single $\varphi$ per electron [72, Sec. 90]. Pseudoscalar and pseudovector operators effectively contribute to those explicitly included in Eq. 8. Consider a background field $\varphi$ with general electron couplings of the form
@@
where again $G_{FF}$ and $G_{F\tilde{F}}$ are scalar functions of $\varphi$ and $\cdots$ designates Lorentz structures with higher derivatives of $\varphi$. The electromagnetic exchange potential (i.e., the amplitude for $ee\rightarrow ee$) receives corrections from all interaction terms, resulting in a coupling function $g_{e}(\varphi)\equiv g_{\alpha}(\varphi)$ that depends on $G_{F\tilde{F}}$ in addition to $G_{FF}$. Therefore, the corrections from derivative couplings are suppressed for the nonrelativistic, ultralight fields we consider here and the scalar coupling fully captures the low-energy physics.
@@ II.2 Effective field theory at low energies
In this section, we discuss the effective field theory description of the interaction between scalars and composite matter. Around the time of primordial nucleosynthesis, the scalar interacts with nucleons and bound nuclei. After recombination, some electrons and nuclei are free while others are bound as neutral atoms. An effective description in terms of these composite states is therefore pertinent to study the scalars interactions with the early-Universe plasma and present-day experimental apparatuses. Here we build upon and expand the discussions in Refs. [22, 21].
@@
The form of the electromagnetic dilatonic charge $(Q_{m_{X}})_{e}$ is less certain due to nuclear effects. Historically, Ref. [20] is the first attempt to estimate $(Q_{m_{X}})_{e}$ in the context of time-varying fundamental constants, arriving at $4(Q_{m_{X}})_{e}\approx 1.3\times 10^{-2}$ by using a classical charge distribution model of the proton [or equivalently by extrapolating the large $Z$ limit of Eq. 21 below to $Z=1$]. More modern methods yield $(Q_{m_{X}})_{e}\sim 10^{-4}$; see Ref. [89, Sec. IIA] for a summary. Refs. [22, 21] estimate $(Q_{m_{X}})_{e}$ using the semi-empirical mass formula for nuclei (SEMF or the Bethe-Weizsäcker formula), as well as the value for the electromagnetic energies of the proton and neutron from [90]. While improvements in nuclear modeling and measurements have further refined these estimates, they nevertheless capture $\mathcal{O}(1)$ of the effects and are sufficient for the present discussion. Therefore, we use Eqs. (16-21) of Ref. [21], extended to ionized species:
@@ II.3.2 Interacting plasma
The nonrelativistic ($T_{p}\ll m_{X}$) and dilute (nondegenerate) limit with a matter-antimatter asymmetry ($n_{X}\gg n_{\bar{X}}$) is $T_{p}\ll m_{X}-\mu_{X}\ll m_{\bar{X}}-\mu_{\bar{X}}$. The interaction contribution is given by [95, see Eq. (5.61)],
@@ II.3.3 After electron-positron annihilation
An additional subtlety is that, after photon decoupling, the electrons and photons do not have a common temperature as assumed in Eq. 34c. Namely, after recombination the temperature of matter redshifts as $1/a^{2}$ rather than $1/a$. However, we only expect the appropriate extension of the bracketed term in Eq. 34c to remain a suppression factor for any temperatures below the electron mass. We therefore ignore this distinction; determining whether the scalar rolls substantially leading up to recombination does not require treating this regime regardless.
In summary, we may conveniently parameterize the thermal potential in terms of effective dilatonic charges for the total baryonic content, i.e., as a weighted sum of those for hydrogen and helium atoms:
@@
Figure 1: Contributions to the effective matter potential Eq. 22 for a scalar coupled to the electron (blue) or photon (red) relative to the total energy density of the Universe. For a quadratically coupled scalar, the vertical axis is equal to $2\delta m_{T}^{2}/3H^{2}d_{\lambda}^{(2)}$ where $\delta m_{T}$ is its in-medium or “thermal” mass. Grey lines depict various mass scales relative to the Hubble rate (both squared) as labeled on the figure; where they intersect the red and blue curves indicates when the in-medium mass takes the labeled value (up to a factor of $23d_{\lambda}^{(2)}/$ 2). Vertical lines indicate when the photon temperature $T=m_{e}$ and $T=T_{\mathrm{eq}}\approx 0.80\leavevmode\nobreak\ \mathrm{eV}$.
@@
where the present-day baryon density $\omega_{b}\equiv\bar{\rho}_{b,0}/3H_{100}^{2}M_{\mathrm{pl}}^{2}$. Its parameterization in terms of the Hubble rate squared in Eq. 38 facilitates identifying under what conditions thermal effects are important to the scalar’s dynamics. Since $\bar{\rho}_{b}/\bar{\rho}_{\mathrm{crit}}\lesssim 1/6$ at all times [99] (with the inequality saturated during the matter-dominated era), matter effects never dominate the scalar’s dynamics after electron-positron annihilation if both quadratic couplings satisfy $d_{\lambda}^{(2)}\lesssim 1/(Q_{b})_{\lambda}\sim 10^{4}$, as also evident in Fig. 1. In Section IV.2.1 we discuss more precise conditions under which even subdominant effects can be neglected within the context of our analysis of variations of fundamental constants as relevant to analyses of cosmological data.
@@ II.5 Cosmological dynamics
where the “thermal mass” $\delta m_{T}$ is given by Eqs. 38 and 41. Note that $\delta m_{T}(a)^{2}$ is proportional to the quadratic couplings $d_{\lambda}^{(2)}$ and therefore is not necessarily positive. The field $\varphi$ itself contributes to the Friedmann equation for the Hubble parameter $H$ but is subdominant so long as $m_{\text{eff}}^{2}\varphi^{2}\ll H^{2}$, which we assume to be the case. To first approximation, expansion thus proceeds as in standard $\Lambda$CDM cosmology.
@@ II.5.1 Free evolution
In the bare-mass–dominated limit ($m_{\varphi}\gg|\delta m_{T}|$), Eq. 45 reduces to the standard equation of motion for a massive scalar, a regime that has been extensively studied in the context where $\varphi$ is, e.g., a pre-inflationary axion, some other light scalar dark matter candidate, or a dynamical subcomponent of dark energy. In such a context, solutions to Eq. 45 are effectively frozen at the initial condition $\bar{\varphi}_{i}$ until $H$ drops below $m_{\varphi}$, at which point damped, harmonic oscillations begin. If the field is sufficiently light at recombination ($m_{\varphi}<H_{\star}\approx 3\times 10^{-29}\leavevmode\nobreak\ \mathrm{eV}$) but is also heavy today ($m_{\varphi}>H_{0}$) oscillations begin between recombination and the present day. The fundamental constants are then spacetime independent over the Universe’s history up to (and including) recombination but different from their current values. Hence, the free evolution of a hyperlight ($10^{-33}\leavevmode\nobreak\ \mathrm{eV}<m_{\varphi}<10^{-28}\leavevmode\nobreak\ \mathrm{eV}$) field realizes the scenario considered in previous phenomenological studies of shifted fundamental constants at BBN and recombination. Here and in Ref. [60] we focus on this “hyperlight” regime.
@@
Solutions in the oscillatory regime may be extended until after matter-$\Lambda$ equality via
@@ III.1 Coherent oscillations
To concretely reason about the effect of phase marginalization, we ground our analysis with a likelihood. To capture the variety of signals we consider, we begin with a general parameterization of a Gaussian likelihood with measurement mean $\hat{x}$ and variance $\sigma_{x}$. We decompose the model prediction at parameters $\bm{\theta}$ into amplitude and phase factors $A(t;\bm{\theta})P(t;\bm{\theta})$, where $P$ is a periodic function bounded by unity in magnitude and $A$ encodes the signal magnitude if $P(t;\bm{\theta})$ were at an antinode. In the relevant limit where the phase $\vartheta$ of $P$ is effectively unknown and treated as a uniform variate, the phase factor simply reduces to $P(t;\bm{\theta})\to P(\vartheta)$. The phase-marginalized posterior is then
@@
with $\pi(\bm{\theta})$ the prior over all other parameters. The $95\%$ highest density interval, used to place upper limits, depends on the behavior of the phase-marginalized posterior in its tails, i.e., where $|A(t;\bm{\theta})|\gg|\hat{x}\pm\sigma_{x}|$. In this regime, the integrand of the marginal posterior is dominated where $|P(\vartheta)|\ll 1$. The only important property of $P(\vartheta)$ is therefore how quickly it approaches zero at its steepest node. If $P(\vartheta)\propto\vartheta^{m}$ (taking the relevant node to be at $\vartheta=0$ for simplicity), the integral’s support is then dominated where $|\vartheta|^{m}<|\hat{x}\pm\sigma_{x}|/A(t;\bm{\theta})$. Approximating the Gaussian as constant over this interval reveals that the marginal posterior scales with $1/\sqrt[m]{A(t;\bm{\theta})}$ when $A(t;\bm{\theta})$ is large, a finding reproduced by numerical quadrature. Given that $A(t;\bm{\theta})$ is often linear in an unknown coupling $d_{\lambda}^{(n)}$, the marginal posterior over the coupling is improper (since the integral over $d_{\lambda}^{(n)}$ does not converge at infinity).
@@ III.2 Spatial variations about a central body
Recall from Eq. 18 that the couplings of composite bulk matter to $\varphi$ are suppressed relative to the fundamental Lagrangian couplings by the dilatonic charges Eq. 21, reflecting that only a fraction of the total energy couples to $\varphi$. The dilatonic charge of an atomic mixture is the average of the dilatonic charges of the atomic constituents, weighted by mass. Earth is approximately $32\%$ iron and $67\%$ silicate $\text{SiO}_{2}$ by mass [106, 107, 108]. Using Eq. 21 and accounting for isotopic abundances [26], one has $(Q_{m_{\text{Fe}}})_{m_{e}}\approx 2.5\times 10^{-4}$ and $(Q_{m_{\text{SiO}_{2}}})_{m_{e}}\approx 2.7\times 10^{-4}$, as well as $(Q_{m_{\text{Fe}}})_{e}\approx 2.6\times 10^{-3}$ and $(Q_{m_{\text{SiO}_{2}}})_{e}\approx 1.6\times 10^{-3}$. All in all, $(Q_{M_{\oplus}})_{m_{e}}\approx 2.6\times 10^{-4}$ and $(Q_{M_{\oplus}})_{e}\approx 1.9\times 10^{-3}$. For the Earth, the screening condition Eq. 66 translates to $d^{(2)}_{M_{\oplus}}\ll 10^{9}$; because the dilatonic charges of Earth are small, screening is only relevant at fundamental couplings much larger than those considered here. In conclusion, the scalar amplitude and time dependence are well approximated by the background density in measurements in the vicinity of the Earth, as relevant for terrestrial probes (Sections III.6 and III.5), while the spatial gradients remain proportional to $d_{M_{\oplus}}^{(2)}$.

### v2

@@ Abstract
Cosmological scalar fields coupled to the Standard Model drive temporal variations in the fundamental constants that grow with redshift, positioning the early Universe as a powerful tool to study such models. We investigate the dynamics and phenomenology of coupled scalars from the early Universe to the present to consistently leverage the myriad searches for time-varying constants and the cosmological signatures of scalars’ gravitational effects. We compute the in-medium contribution from Standard Model particles to the scalar’s dynamics and identify only a limited range of couplings for which the scalar has an observable impact on the fundamental constants without either evolving before recombination or gravitating non-negligibly. We then extend existing laboratory and astrophysical bounds to the hyperlight scalar regime. We present joint limits from the early and late Universe, specializing to hyperlight, quadratically coupled scalars that modulate the mass of the electron or the strength of electromagnetism and make up a subcomponent of the dark matter today. Our dedicated analysis of observations of the cosmic microwave background, baryon acoustic oscillations, and type Ia supernovae provides the most stringent constraints on quadratically coupled scalars with masses from $10^{-28.5}$ to $\sim 10^{-31}\leavevmode\nobreak\ \mathrm{eV}$, below which quasar absorption spectra yield stronger bounds. These results jointly limit hyperlight scalars that comprise a few percent of the current dark matter density to near- or subgravitational couplings to electrons or photons.
@@ I Introduction
Scalar fields have cosmological consequences as well—even simply through gravity if they are sufficiently abundant. A cosmological density of the new scalar may be produced in the very early Universe, for instance through the misalignment mechanism [29, 30, 31, 32, 33, 13, 34, 35, 36]. Depending on its initial conditions and interactions, the scalar makes up at least a subcomponent of the cosmological dark sector today as matter, radiation, or something more exotic. In general, scalar fields redshift with the expansion of the Universe and their observational effects are more pronounced earlier in cosmological history—more so if the leading dependence of fundamental constants on the scalar field amplitude is quadratic [23, 26, 37, 38] or higher rather than linear. In an era of precision cosmology, early-Universe observations are particularly apt probes of dynamical scalar fields and may explore parameter space complementary to that from traditional particle physics tests.
@@
We use natural units in which $\hbar=c=1$ and define the reduced Planck mass $M_{\mathrm{pl}}\equiv 1/\sqrt{8\pi G}=2.435\times 10^{18}\leavevmode\nobreak\ \mathrm{GeV}$. We employ the Einstein summation convention for spacetime indices; repeated spatial indices (Latin characters) are implicitly summed regardless of their placement. We also use upright boldface to denote spatial vectors. Unless otherwise specified, we fix a homogeneous Friedmann-Lemaître-Robertson-Walker spacetime with metric
@@ II.1 Varying constants with new scalars
The assumed suppression of terms linear in $\varphi$ could be achieved with a $\mathbb{Z}^{\varphi}_{2}:\varphi\mapsto-\varphi$ symmetry; we mention models that feature such a symmetry in the following.
@@
Depending on the assumed symmetries, scalars may interact with fermions and gauge bosons via operators other than those parametrized by Eq. 8. For instance, a shift-symmetric pseudovector coupling is characteristic of (pseudo)Nambu-Goldstone bosons such as axionlike particles [68, 69, 70, 71]. A pseudoscalar coupling is also allowed by symmetries and is equivalent to the pseudovector coupling alone only in processes whose Feynman diagrams have a single $\varphi$ per electron [72, Sec. 90]. Pseudoscalar (PS) and pseudovector (PV) operators effectively contribute to those explicitly included in Eq. 8. Consider a background field $\varphi$ with general electron couplings of the form
@@
where again $G_{FF}$ and $G_{F\tilde{F}}$ are scalar functions of $\varphi$ and $\cdots$ designates Lorentz structures with higher derivatives of $\varphi$. The phenomenology of the latter coupling has been studied extensively and for instance can source cosmic birefringence [73, 74, 75, 76, 77, 78, 79]. Models can also simultaneously feature both couplings, which transform differently under parity [80, 81, 82]. For our purposes, the electromagnetic exchange potential (i.e., the amplitude for $ee\rightarrow ee$) receives corrections from all interaction terms, resulting in a coupling function $g_{e}(\varphi)\equiv g_{\alpha}(\varphi)$ that depends on $G_{F\tilde{F}}$ in addition to $G_{FF}$. However, $F_{\mu\nu}\tilde{F}^{\mu\nu}$ by itself is a total derivative term, which holds for $G_{F\tilde{F}}(\varphi)F_{\mu\nu}\tilde{F}^{\mu\nu}$ unless $\varphi$ varies in spacetime. For an Abelian gauge boson like the photon, any physical effect must therefore be proportional to derivatives of $\varphi$.
Thus, both the PV coupling in Eq. 12 and the $F\tilde{F}$ coupling in Eq. 14 are “derivative” couplings, the corrections from which are suppressed for the nonrelativistic, ultralight fields we consider here and the non-derivative, scalar couplings fully captures the low-energy physics.
@@ II.2 Effective field theory at low energies
In this section, we discuss the effective field theory description of the interaction between scalars and composite matter. Around the time of primordial nucleosynthesis, the scalar interacts with nucleons and bound nuclei. After recombination, some electrons and nuclei are free while others are bound as neutral atoms. An effective description in terms of these composite states is therefore pertinent to study the scalar’s interactions with the early Universe plasma and present-day experimental apparatuses. Here we build upon and expand the discussions in Refs. [22, 21].
@@
The form of the electromagnetic dilatonic charge $(Q_{m_{X}})_{e}$ is less certain due to nuclear effects. Historically, Ref. [20] is the first attempt to estimate $(Q_{m_{X}})_{e}$ in the context of time-varying fundamental constants, arriving at $4(Q_{m_{X}})_{e}\approx 1.3\times 10^{-2}$ by using a classical charge distribution model of the proton [or equivalently by extrapolating the large $Z$ limit of Eq. 21 below to $Z=1$]. More modern methods yield $(Q_{m_{X}})_{e}\sim 10^{-4}$; see Ref. [99, Sec. IIA] for a summary. References [22, 21] estimate $(Q_{m_{X}})_{e}$ using the semi-empirical mass formula for nuclei (SEMF or the Bethe-Weizsäcker formula), as well as the value for the electromagnetic energies of the proton and neutron from [100]. While improvements in nuclear modeling and measurements have further refined these estimates, they nevertheless capture $\mathcal{O}(1)$ of the effects and are sufficient for the present discussion. Therefore, we use Eqs. (16-21) of Ref. [21], extended to ionized species:
@@ II.3.2 Interacting plasma
The nonrelativistic ($T_{p}\ll m_{X}$) and dilute (nondegenerate) limit with a matter-antimatter asymmetry ($n_{X}\gg n_{\bar{X}}$) is $T_{p}\ll m_{X}-\mu_{X}\ll m_{\bar{X}}-\mu_{\bar{X}}$. The interaction contribution is given by [104, see Eq. (5.61)],
@@ II.3.3 After electron-positron annihilation
[…]
In summary, we may conveniently parametrize the thermal potential in terms of effective dilatonic charges for the total baryonic content, i.e., as a weighted sum of those for hydrogen and helium atoms:
@@
Figure 1: Contributions to the effective matter potential Eq. 22 for a scalar coupled to the electron (blue) or photon (red) relative to the total energy density of the Universe. For a quadratically coupled scalar, the vertical axis is equal to $2\delta m_{T}^{2}/3H^{2}d_{\lambda}^{(2)}$ where $\delta m_{T}$ is its in-medium or “thermal” mass. Grey lines depict various mass scales relative to the Hubble rate (both squared) as labeled on the figure; where they intersect the red and blue curves indicates when the in-medium mass takes the labeled value (up to a factor of $3d_{\lambda}^{(2)}/2$). Vertical lines indicate when the photon temperature $T=m_{e}$ and $T=T_{\mathrm{eq}}\approx 0.80\leavevmode\nobreak\ \mathrm{eV}$.
@@
where the present-day baryon density $\omega_{b}\equiv\bar{\rho}_{b,0}/3H_{100}^{2}M_{\mathrm{pl}}^{2}$. Its parametrization in terms of the Hubble rate squared in Eq. 38 facilitates identifying under what conditions thermal effects are important to the scalar’s dynamics. Since $\bar{\rho}_{b}/\bar{\rho}_{\mathrm{crit}}\lesssim 1/6$ at all times [108] (with the inequality saturated during the matter-dominated era), matter effects never dominate the scalar’s dynamics after electron-positron annihilation if both quadratic couplings satisfy $d_{\lambda}^{(2)}\lesssim 1/(Q_{b})_{\lambda}\sim 10^{4}$, as also evident in Fig. 1. In Section IV.2.1 we discuss more precise conditions under which even subdominant effects can be neglected within the context of our analysis of variations of fundamental constants as relevant to analyses of cosmological data.
@@ II.5 Cosmological dynamics
where the “thermal mass” $\delta m_{T}$ is given by Eqs. 38 and 41. Note that $\delta m_{T}(a)^{2}$ is proportional to the quadratic couplings $d_{\lambda}^{(2)}$ and therefore is not necessarily positive. The field $\varphi$ itself contributes to the Friedmann equation for the Hubble parameter $H$ but is subdominant so long as $m_{\text{eff}}^{2}\varphi^{2}\ll H^{2}$, which we assume to be the case. To first approximation, expansion thus proceeds as in standard $\Lambda$ cold dark matter ($\Lambda$CDM) cosmology.
@@ II.5.1 Free evolution
In the bare-mass–dominated limit ($m_{\varphi}\gg|\delta m_{T}|$), Eq. 45 reduces to the standard equation of motion for a massive scalar, a regime that has been extensively studied in the context where $\varphi$ is, e.g., a pre-inflationary axion, some other light scalar dark matter candidate, or a dynamical subcomponent of dark energy. In such a context, solutions to Eq. 45 are effectively frozen at the initial condition $\bar{\varphi}_{i}$ until $H$ drops below $m_{\varphi}$, at which point damped, harmonic oscillations begin. If the field is sufficiently light at recombination ($m_{\varphi}<H_{\star}\approx 3\times 10^{-29}\leavevmode\nobreak\ \mathrm{eV}$) but is also heavy today ($m_{\varphi}>H_{0}$) oscillations begin between recombination and the present day. The fundamental constants are then spacetime independent over the Universe’s history up to (and including) recombination but different from their current values. Hence, the free evolution of a hyperlight ($10^{-33}\leavevmode\nobreak\ \mathrm{eV}<m_{\varphi}<10^{-28}\leavevmode\nobreak\ \mathrm{eV}$) field realizes the scenario considered in previous phenomenological studies of shifted fundamental constants at big bang nucleosynthesis (BBN) and recombination. Here and in Ref. [60] we focus on this “hyperlight” regime.
@@
[…]
@@ III.1 Coherent oscillations
To concretely reason about the effect of phase marginalization, we ground our analysis with a likelihood. To capture the variety of signals we consider, we begin with a general parametrization of a Gaussian likelihood with measurement mean $\hat{x}$ and variance $\sigma_{x}$. We decompose the model prediction at parameters $\bm{\theta}$ into amplitude and phase factors $A(t;\bm{\theta})P(t;\bm{\theta})$, where $P$ is a periodic function bounded by unity in magnitude and $A$ encodes the signal magnitude if $P(t;\bm{\theta})$ were at an antinode. In the relevant limit where the phase $\vartheta$ of $P$ is effectively unknown and treated as a uniform variate, the phase factor simply reduces to $P(t;\bm{\theta})\to P(\vartheta)$. The phase-marginalized posterior is then
@@
[…]
@@
Finally, we reiterate that phase marginalization does not so severely penalize the actual discovery potential of a given search: while a null signal can be consistent with a large coupling if the scalar were near a node, a measurement of a nonzero signal cannot be explained by an arbitrarily small coupling, no matter the scalar’s phase. The above arguments concern measures of the tail of posterior distributions and still apply in that regime, but the posterior would still localize about finite couplings if the model could match the detected signal. When observing only a small fraction of an oscillation, however, it may be challenging to disentangle the phase factor $P(\vartheta)$ from the actual amplitude $A(t;\bm{\theta})$ and measure the latter at better than the order-unity level.
@@ III.2 Spatial variations about a central body
[…]